# DBgDel: Database-Enhanced Gene Deletion Framework for Growth-Coupled Production in Genome-Scale Metabolic Models

### v1

@@ I INTRODUCTION
During simulation in constraint-based models, the primary objective is to achieve growth-coupled production for target metabolites. Growth-coupled production means cell growth is bonded to the synthesis of target metabolites in the metabolic process of microorganisms. Coupling cell growth with the production of target metabolites is necessary. The reason is in industrial practices, microorganism genotypes with higher cell growth are more likely to persist in the culture through repeated passaging. However, in the natural metabolic state of most microorganisms, only a limited number of metabolites meet the criteria for growth-coupled production. Therefore, achieving growth-coupled production for most target metabolites requires calculating gene deletion strategies to reshape the metabolic network [6]. But this task is far from trivial. Calculating gene deletion strategies demands huge computational resources, especially when simulating genome-scale models with complex metabolic networks and GPR rules involving many genes.
@@
Several methods have been proposed to address this issue. Among the existing approaches [7, 8, 9, 10, 11, 12, 13, 14] for calculating gene deletion strategies in growth-coupled production, one of the most efficient is the elementary flux vector-based method [15]. Its basic idea is to identify a minimal set of reactions in the flow where cell growth compels the production of the target metabolite. Specifically, the elementary flux vector-based method determines a non-decomposable flux distribution that encompasses (1) the cell growth reaction and (2) the target metabolite production reaction, then the reactions not utilized by the flux distribution are eliminated. This method has demonstrated its capability to compute reaction deletions for growth-coupled production for most target metabolites in common model microorganisms like E. coli and S. cerevisiae [15]. Furthermore, the elementary flux vector-based method has been successfully extended and applied to calculate the growth-coupled production of valuable metabolites, including itaconic acid and 2,3-butanediol in E. coli [16], as well as indigoidine in P. putida [17]. More recently, Tamura et al. introduced gDel_minRN [18], a tool to calculate gene deletion strategies. It operates by maximizing the number of repressed reactions, thereby extracting essential core components for growth-coupled production. Experimental validations have demonstrated that in comparison with other methods, gDel_minRN stands out as one of the most effective approaches currently available for calculating gene deletions. However, despite numerous efforts, these approaches exhibit a common shortcoming. All of the present methods employ de novo calculation strategies when deriving gene deletion strategies for various target metabolites. This approach overlooks the shared information among different target metabolites, potentially leading to an excessive number of unnecessary repeated calculations.
@@
On the other side, efforts have been made to establish databases on gene deletion strategies to address the growing demand for computational experiment data in strain design. A noteworthy example is the MetNetComp [19] database. It is a web-based system that offers information on gene deletion strategies for growth-coupled production in constraint-based metabolic networks. MetNetComp encompasses a vast repository of over 85,000 gene deletion strategies for metabolites across various constraint-based models from different species. While the advent of such databases has significant potential to foster new research paradigms in the field, the exploration of their utility, especially concerning computational gene deletion strategies, is still in its infancy.
@@
In this study, the authors propose DBgDel, a novel database-enhance framework for calculating gene deletion strategies for growth-coupled production on genome-scale models. DBgDel extracts information on gene deletion from the related database to boost the computational efficiency of downstream algorithms. DBgDel comprises two steps: (1) STEP 1, extracting the gene set from a gene deletion database that corresponds to the essential core components for growth-coupled production for a given constraint-based model; and (2) STEP 2, using an extended version of gDel_minRN that incorporates the gene set extracted from the STEP 1 as the initial gene pool to narrow down the algorithmic search space.
In the computational experiments, we compared DBgDel with gDel_minRN [18], GDLS [7], and optGene [20]. Besides the aforementioned algorithm gDel_minRN, GDLS and optGene are also the most widely used software to derive gene deletion strategies, which are available in the COBRA Toolbox [21]. All these methods were applied to three constraint-based models including e_coli_core [22], iMM904 [23], and iML1515 [24]. e_coli_core is a constraint-based model that contains the only essential part of the metabolism of E. coli; iMM904 and iML1515 are genome-scale constraint-based models of S. cerevisiae and E. coli, respectively;
The average elapsed time of DBgDel for e_coli_core, iMM904, and iML1515 were 1.12s, 79.92s, and 431.75s, respectively, which were substantially faster than gDel_minRN, GDLS and optGene. In the meantime, DBgDel achieves the success ratio for e_coli_core, iMM904, and iML1515 at 60.0%, 11.5%, and 51.2%, which were substantially better than GDLS and optGene, and closely similar to gDel_minRN .
@@
The remaining sections of this paper are organized as follows: Section II-A formularizes several fundamental concepts in this study; Section II-B and II-C describes the main problem of this study mathematically and illustrates it with a small example; Section III illustrates the proposed framework DBgDel with a small example and provides the corresponding pseudo-code; Section IV-A describes the basic experiment setting; Section IV-B describes the experiment results: (1) the performance comparison of DBgDel, gDel_minRN, GDLS, and optGene for e_coli_core, iMM904, and iML1515, (2) the performance comparison of DBgDel based on different initial gene pools, including Predicted-$G_{\text{remain}}$ genes as the default setting, $G_{\text{remain}}$ genes, and growth essential genes; Section V analyzes the results of the experiments, evaluates the performance of DBgDel and other methods, and discusses future work.
@@ II-A2 Flux balance analysis
When analyzing the metabolic network within a constraint-based model, flux balance analysis (FBA) assumes steady states where all metabolic reaction speeds (fluxes) are constant [25]. Specifically, we give the following definitions for FBA: (1) for each compound, the sum of the producing fluxes is equal to the sum of the consuming fluxes; (2) in each reaction, the fluxes of substrates and products must satisfy the ratio in the chemical reaction equation, and (3) the upper and lower bounds are given for each flux. In the standard procedure for FBA on the given metabolic network $C_{1}$, the objective is to maximize the cell growth reaction rate $v_{\text{growth}}$ using the following linear programming (LP):
@@ II-C Example
Figure 1: A toy example of the constraint-based model where circles and rectangles represent metabolites and reactions, respectively. Black rectangles denote external and internal reactions. $r_{1}$, $r_{2}$ correspond to two substrate uptake reactions. $r_{8}$, $r_{9}$ correspond to cell growth, and target metabolite production reactions, respectively. The reaction speeds are constrained by the range $[l_{i},u_{i}]$. This example shows only part of the model, the rest after $m_{7}$ is omitted.
@@ III-A2 STEP 2
29:            s.t. max $v_{\text{growth}}$
30:            s.t. $\sum_{j}S_{ij}v_{j}=0$ for all $i$
31:            $\begin{cases}v_{j}=0&\text{ if }p_{j}=0,\\ l_{j}\leq v_{j}\leq u_{j},&\text{otherwise};\end{cases}$
32:            $p_{j}=f_{j}(G)$
33:            $g=\begin{cases}0&\text{if }g\in D_{\text{candidate}},\\ 1,&\text{otherwise};\end{cases}$
34:     if $v_{\text{target}}\geq PR_{\text{threshold}}$ and $v_{\text{growth}}\geq GR_{\text{threshold}}$ then
35:        return  $D_{\text{candidate}}$, $v_{\text{target}}$, $v_{\text{growth}}$
36:     else
37:        $prohibited\_list=prohibited\_list\cup D_{\text{candidate}}$
38:        $loop=loop+1$
39:     end if
40:  end while
@@ III-B Example
Figure 3: The toy model with a new target metabolite. Circles and rectangles represent metabolites and reactions, respectively. Black rectangles denote external and internal reactions. $r_{1}$, $r_{2}$ correspond to two substrate uptake reactions. $r_{8}$, $r_{10}$ correspond to cell growth, and target metabolite production reactions, respectively. The reaction speeds are constrained by the range $[l_{i},u_{i}]$.
@@ IV-A Experiments Setting
We define three types of genes used in the experiments, i.e., growth essential (GE) genes, $G_{\text{remain}}$ genes, and Predicted-$G_{\text{remain}}$ genes as below:
@@
    Evaluate the overall computational efficiency of different gene deletion calculation methods. We compared the performance of DBgDel with other methods, including gDel_minRN, GDLS, and optGene.
@@
    Evaluate the efficacy of different initial gene pool settings of DBgDel . We compared the performance of DBgDel with different initial gene pools in STEP2: Predicted-$G_{\text{remain}}$ genes, GE genes, and $G_{\text{remain}}$ genes. Here, Predicted-$G_{\text{remain}}$ is the default choice of the initial gene pool, and $G_{\text{remain}}$ serves as the performance benchmark.
@@
All procedures in the computational experiments were implemented on a Ubuntu 20.04.6 LTS machine with an Intel Xeon Gold Processor with 2.30 GHz 64 cores/128 threads, 128 GB memory, and 1TB SSD. This workstation had CPLEX 12.10, COBRA Toolbox v3.0, and MATLAB R2019b installed and used for these analyses.
@@ IV-B Results and Performance Comparison
For e_coli_core: Table VIII summarizes the performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, and gDel_minRN in the context of targeting 40 non-extracellular space metabolites of e_coli_core. Regarding the success rate, DBgDel achieved success in 24 out of 40 target metabolites, meeting the minimum GR and PR threshold of 0.001 or higher during GR maximization. In comparison, GDLS succeeded in only 5 out of 40 cases, optGene in 22 out of 40, and gDel_minRN in 28 out of 40 cases. Regarding computational efficiency, optGene required significantly more extended time, with an average of 1779.30 seconds, while the other three methods cost about 1 second.
Table IX summarizes the performance comparison between the proposed method DBgDel with two different initial remaining gene pool settings. When utilizing Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate in achieving the minimum GR and PR threshold of 0.001 or higher during GR maximization was 24 out of 40 instances. Similarly, when using the GE gene as the initial remaining gene pool, the success rate was 28 out of 40 instances. Additionally, when employing $G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate remained at 28 out of 40 instances. As for computational efficiency, since the model scale of e_coli_core is relatively small, all three settings of the initial remaining gene pool can achieve a computation time of around 1 second, which is short enough.
TABLE VIII: The performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, and gDel_minRN in the situation of targeting non-extracellular space metabolites of e_coli_core. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GDLS optGene gDel_minRN
Success rate 24/40 5/40 22/40 28/40
Avg.elapsed time 1.12 1.24 1779.30 1.62
TABLE IX: The performance comparison between the proposed method DBgDel with two different initial remaining gene pool settings for e_coli_core. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GE gene $G_{\text{remain}}$
Success rate 24/40 28/40 28/40
Time.Success = 1 0.59 0.68 0.60
Time.Success = 0 1.91 2.83 2.57
Time.All cases 1.12 1.33 1.12
@@
For iMM904: Table X summarizes the performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, and gDel_minRN in the context of targeting 700 non-extracellular space metabolites of iMM904. Regarding the success rate, DBgDel achieved success in 81 out of 700 target metabolites, meeting the minimum GR and PR threshold of 0.001 or higher during GR maximization. In comparison, GDLS failed in all 700 cases, optGene succeeded in only 27 out of 700 cases, and gDel_minRN in 116 out of 700 cases. DBgDel demonstrated outstanding performance in average elapsed time, with a mean time of 79.79 seconds. GDLS had an average time of 243.42 seconds, optGene required significantly more extended time with an average of 1986.45 seconds, and gDel_minRN had an average time of 422.32 seconds.
@@
Table XI summarizes the performance comparison between the proposed method DBgDel with two different initial remaining gene pool settings. When utilizing Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate in achieving the minimum GR and PR threshold of 0.001 or higher during GR maximization was 81 out of 700 instances. Similarly, when using the GE gene as the initial remaining gene pool, the success rate was 120 out of 700 instances. Additionally, when employing $G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate remained at 126 out of 700 instances. Examining the computational efficiency, the proposed method demonstrated varied performance regarding elapsed time for successful cases. For instances where success was achieved, the average elapsed time was 22.89 seconds with Predicted-$G_{\text{remain}}$, 189.90 seconds with GE gene, and 47.82 seconds with $G_{\text{remain}}$ genes. In contrast, for cases where success was not achieved, the average elapsed time was 87.38 seconds with Predicted-$G_{\text{remain}}$, 459.31 seconds with GE gene, and 224.19 seconds with $G_{\text{remain}}$ genes. Considering all cases, regardless of success or failure, the proposed method showed outstanding average elapsed times when using Predicted-$G_{\text{remain}}$ or $G_{\text{remain}}$ genes, with values of 79.92 and 192.44 seconds. The average elapsed time of 413.13 seconds is still competitive when using the GE gene.
@@
TABLE X: The performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, and gDel_minRN in the situation of targeting non-extracellular space metabolites of iMM904. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GDLS optGene gDel_minRN
Success rate 81/700 0/700 27/700 116/700
Avg.elapsed time 79.92 243.42 1986.45 422.32
TABLE XI: The performance comparison between the proposed method DBgDel with two different initial remaining gene pool settings for iMM904. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GE gene $G_{\text{remain}}$
Success rate 81/700 120/700 126/700
Time.Success = 1 22.89 189.90 47.82
Time.Success = 0 87.38 459.31 224.19
Time.All cases 79.92 413.13 192.44
@@
For iML1515: Table XII summarizes the performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, and gDel_minRN in the context of targeting 991 non-extracellular space metabolites of iML1515. Regarding the success rate, DBgDel achieved success in 507 out of 991 target metabolites, meeting the minimum GR and PR threshold of 0.001 or higher during GR maximization. GDLS and optGene failed in all 991 cases, and gDel_minRN succeeded in 508 out of 991 cases. DBgDel demonstrated outstanding performance with an average elapsed time of 431.75 seconds. GDLS had an average time of 244.58 seconds, optGene required significantly more extended time with an average of 2193.45 seconds, and gDel_minRN had an average time of 2541.23 seconds.
@@
Table XIII summarizes the performance comparison between the proposed method DBgDel with two different initial remaining gene pool settings. When utilizing Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate in achieving the minimum GR and PR threshold of 0.001 or higher during GR maximization was 507 out of 991 instances. Similarly, when using the GE gene as the initial remaining gene pool, the success rate was 508 out of 991 instances. Additionally, when employing $G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate remained at 508 out of 991 instances. Examining the computational efficiency, the proposed method demonstrated varied performance regarding elapsed time for successful cases. For instances where success was achieved, the average elapsed time was 92.46 seconds with Predicted-$G_{\text{remain}}$, 481.66 seconds with GE gene, and 90.30 seconds with $G_{\text{remain}}$ genes. In contrast, for cases where success was not achieved, the average elapsed time was 787.17 seconds with Predicted-$G_{\text{remain}}$, 4561.56 seconds with GE gene, and 858.01 seconds with $G_{\text{remain}}$ genes. Considering all cases, regardless of success or failure, the proposed method showed outstanding average elapsed times when using Predicted-$G_{\text{remain}}$ or $G_{\text{remain}}$ genes, with values of 79.92 and 192.44 seconds. The average elapsed time of 2470.15 seconds is still competitive when using the GE gene.

### v3

@@ I INTRODUCTION
During the simulation of gene deletion in constraint-based models, the primary objective is to achieve growth-coupled production for target metabolites. Growth-coupled production means cell growth is bonded to the synthesis of target metabolites in the metabolic process of microorganisms. Specifically, in this study, we adopt the paradigm of weakly growth-coupled production, which means achieving a non-zero synthesis rate of the target metabolite at the maximum non-zero growth rate [6, 7]. Coupling cell growth with the production of target metabolites is necessary. The reason is in industrial practices, microorganism genotypes with higher cell growth are more likely to persist in the culture through repeated passaging. However, in the natural metabolic state of most microorganisms, only a limited number of metabolites meet the criteria for growth-coupled production. Therefore, achieving growth-coupled production for most target metabolites requires calculating gene deletion strategies to reshape the metabolic network [8]. However, gene deletion strategies calculation tasks are far from trivial. Calculating gene deletion strategies demands huge computational resources, especially when simulating genome-scale models with complex metabolic networks and GPR rules involving many genes.
@@
Several methods have been proposed to address this issue. Among the existing approaches [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] for calculating gene deletion strategies in growth-coupled production, the most efficient ones include the minimal cut set (MCS)-based method that originated from the elementary flux vector (EFV)-based method [20]. The basic idea of the EFV method is to identify a minimal set of reactions in the flow where cell growth compels the production of the target metabolite. Specifically, it determines a non-decomposable flux distribution that encompasses (1) the cell growth reaction and (2) the target metabolite production reaction, then the reactions not utilized by the flux distribution are eliminated. The MCS method identifies the smallest possible set of gene knockouts, with computations based on Farkas Lemma [21]. MCS method has demonstrated its capability to compute reaction deletions for growth-coupled production for most target metabolites in common model microorganisms like E. coli and S. cerevisiae [20]. Furthermore, these methods have been successfully extended and applied to calculate the growth-coupled production of valuable metabolites, including itaconic acid and 2,3-butanediol in E. coli [22], as well as indigoidine in P. putida [23]. More recently, Tamura et al. introduced gDel_minRN [24], a tool to calculate gene deletion strategies. It operates by maximizing the number of repressed reactions, thereby extracting essential core components for growth-coupled production. Experimental validations have demonstrated that in comparison with other methods, gDel_minRN stands out as one of the most effective approaches currently available for calculating gene deletions. However, despite numerous efforts, these approaches exhibit a common shortcoming. All of the present methods employ de novo calculation strategies when deriving gene deletion strategies for various target metabolites. This approach overlooks the shared information among different target metabolites, potentially leading to an excessive number of unnecessary repeated calculations.
@@
On the other side, efforts have been made to establish databases on gene deletion strategies to address the growing demand for computational experiment data in strain design. A noteworthy example is the MetNetComp [25] database. It is a web-based system that offers information on gene deletion strategies for growth-coupled production in constraint-based metabolic networks. MetNetComp computed growth-coupled gene deletion strategies that are minimal or maximal regarding the number of gene deletions by modifying the gDel_minRN strategies. It now houses a vast repository of over 85,000 gene-deletion strategies for metabolites across various constraint-based models from different species. While the advent of such databases has significant potential to foster new research paradigms in the field, the exploration of their utility, especially concerning computational gene deletion strategies, is still in its infancy.
@@
In this study, the authors propose DBgDel, a novel database-enhance framework for calculating gene deletion strategies for growth-coupled production on genome-scale models. DBgDel extracts information on gene deletion from the related database to boost the computational efficiency of downstream algorithms. DBgDel comprises two steps: (1) STEP 1, extracting the gene set from a gene deletion database that corresponds to the essential core components for growth-coupled production for a given constraint-based model; and (2) STEP 2, using an extended version of gDel_minRN that incorporates the gene set extracted from the STEP 1 as the initial gene pool. The initial gene pool contains genes that are never knocked out in the gene deletion database, and these are protected from knockouts to narrow the algorithmic search space.
In the computational experiments, we compared DBgDel with GDLS [12], optGene [26], gDel_minRN [24], and geneMCSEnumerator (gMCSE) [20]. Besides the aforementioned algorithm gDel_minRN, gMCSE is an MCS-based algorithm for growth-coupled strain design using gene knockouts, which is available as an API function ’geneMCSEnumerator2’ in the CellNetAnalyzer [27], and we evaluated its first resulting mcs. GDLS and optGene are also the most widely used software to derive gene deletion strategies, which are available in the COBRA Toolbox [28].
All these methods were applied to three constraint-based models including e_coli_core [29], iMM904 [30], and iML1515 [31]. e_coli_core is a constraint-based model that contains the only essential part of the metabolism of E. coli; iMM904 and iML1515 are genome-scale constraint-based models of S. cerevisiae and E. coli, respectively;
The average elapsed time of DBgDel for e_coli_core, iMM904, and iML1515 were 1.12s, 79.92s, and 431.75s, respectively, which were substantially faster than all other compared methods. In the meantime, DBgDel achieves the success ratio for e_coli_core, iMM904, and iML1515 at 60.0%, 11.5%, and 51.2%, which were substantially better than GDLS, optGene, and gMCSE, and closely similar to gDel_minRN.
@@
The remaining sections of this paper are organized as follows: Section II-A formularizes several fundamental concepts in this study; Section II-B and II-C describes the main problem of this study mathematically and illustrates it with a small example; Section III illustrates the proposed framework DBgDel with a small example and provides the corresponding pseudo-code; Section IV-A describes the basic experiment setting; Section IV-B describes the experiment results: (1) the performance comparison of DBgDel, GDLS, optGene, gDel_minRN, and gMCSE for e_coli_core, iMM904, and iML1515, (2) the performance comparison of DBgDel based on different initial gene pools, including Predicted-$G_{\text{remain}}$ genes as the default setting, $G_{\text{remain}}$ genes, growth essential genes, and randomly chosen genes; Section V analyzes the results of the experiments, evaluates the performance of DBgDel and other methods, and discusses future work.
@@ II-A2 Flux balance analysis
When analyzing the metabolic network within a constraint-based model, flux balance analysis (FBA) assumes steady states where all metabolic reaction rates (fluxes) are constant [32]. Specifically, we give the following definitions for FBA: (1) for each compound, the sum of the producing fluxes is equal to the sum of the consuming fluxes; (2) in each reaction, the fluxes of substrates and products must satisfy the ratio in the chemical reaction equation, and (3) the upper and lower bounds are given for each flux. In the standard procedure for FBA on the given metabolic network $C_{1}$, the objective is to maximize the cell growth reaction rate $v_{\text{growth}}$ using the following linear programming (LP):
@@ II-C Example
Figure 1: A toy example of the constraint-based model where circles and rectangles represent metabolites and reactions, respectively. Black rectangles denote external and internal reactions. $r_{1}$, $r_{2}$ correspond to two substrate uptake reactions. $r_{8}$, $r_{9}$ correspond to cell growth, and target metabolite production reactions, respectively. The reaction rates are constrained by the range $[l_{i},u_{i}]$. This example shows only part of the model, the rest after $m_{7}$ is omitted.
@@ III-A2 STEP 2
29:         such that max $v_{\text{growth}}$
30:            such that
31:               $\sum_{j}S_{ij}v_{j}=0$ for all $i$
32:               $\begin{cases}v_{j}=0&\text{ if }p_{j}=0,\\ l_{j}\leq v_{j}\leq u_{j},&\text{otherwise};\end{cases}$
33:               $p_{j}=f_{j}(G)$
34:               $g=\begin{cases}0&\text{if }g\in D_{\text{candidate}},\\ 1,&\text{otherwise};\end{cases}$
35:     if $v_{\text{target}}\geq PR_{\text{threshold}}$ and $v_{\text{growth}}\geq GR_{\text{threshold}}$ then
36:        return  $D_{\text{candidate}}$, $v_{\text{target}}$, $v_{\text{growth}}$
37:     else
38:        $prohibited\_list=prohibited\_list\cup D_{\text{candidate}}$
39:        $loop=loop+1$
40:     end if
41:  end while
@@ III-B Example
Figure 3: The toy model with a new target metabolite. Circles and rectangles represent metabolites and reactions, respectively. Black rectangles denote external and internal reactions. $r_{1}$, $r_{2}$ correspond to two substrate uptake reactions. $r_{8}$, $r_{10}$ correspond to cell growth, and target metabolite production reactions, respectively. The reaction rates are constrained by the range $[l_{i},u_{i}]$.
@@ IV-A Experiments Setting
We define four types of genes used in the experiments, i.e., Growth Essential (GE) genes, $G_{\text{remain}}$ genes, Predicted-$G_{\text{remain}}$ genes, and Randomly Chosen (RC) genes as below:
@@
  • Randomly Chosen (RC) genes: A set of genes randomly chosen from the given metabolic model. RC genes have no biological significance and are used to ensure a fair comparison of effectiveness in narrowing the search space. The number of RC genes equals that of Predicted-$G_{\text{remain}}$ genes.
@@
    Evaluate the overall computational efficiency of different gene deletion calculation methods. We compared the performance of DBgDel with other baseline methods, including GDLS, optGene, gDel_minRN, and gMCSE.
@@
    Evaluate the efficacy of different initial gene pool settings of DBgDel. We compared the performance of DBgDel with different initial gene pools in STEP2: Predicted-$G_{\text{remain}}$ genes, GE genes, $G_{\text{remain}}$ genes, and RC genes. Here, Predicted-$G_{\text{remain}}$ is the default choice of the initial gene pool, and $G_{\text{remain}}$ serves as the performance benchmark.
@@
All procedures in the computational experiments were implemented on a Ubuntu 20.04.6 LTS machine with an Intel Xeon Gold Processor with 2.30 GHz 64 cores/128 threads, 128 GB memory, and 1TB SSD. This workstation had CPLEX 12.10, COBRA Toolbox v3.0, CellNetAnalyzer ver.2023.1, and MATLAB R2019b installed and used for these analyses.
@@ IV-B Results and Performance Comparison
For e_coli_core: Table VIII summarizes the performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, gDel_minRN, and gMCSE in the context of targeting 40 non-extracellular space metabolites of e_coli_core. Regarding the success rate, DBgDel achieved success in 24 out of 40 target metabolites, meeting the minimum GR and PR threshold of 0.001 or higher during GR maximization. In comparison, GDLS succeeded in only 5 out of 40 cases, optGene in 22 out of 40 cases, gDel_minRN in 28 out of 40 cases, and gMCSE in 7 out of 40 cases. Regarding computational efficiency, optGene required significantly more extended time, with an average of 1779.30 seconds, while the other methods cost about 1 second.
Table IX summarizes the performance comparison between the proposed method DBgDel with three different initial remaining gene pool settings. When utilizing Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate in achieving the minimum GR and PR threshold of 0.001 or higher during GR maximization was 24 out of 40 instances. Similarly, when using the GE genes as the initial remaining gene pool, the success rate was 28 out of 40 instances. When employing $G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate remained at 28 out of 40 instances. Additionally, when employing RC genes as the initial remaining gene pool, the success rate decreased to 21 out of 40 instances. As for computational efficiency, since the model scale of e_coli_core is relatively small, all four settings of the initial remaining gene pool can achieve a computation time of around 1 second, which is short enough.
TABLE VIII: The performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, gDel_minRN, and gMCSE in the situation of targeting non-extracellular space metabolites of e_coli_core. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GDLS gMCSE optGene gDel_minRN
Success rate 24/40 5/40 7/40 22/40 28/40
Avg. time 1.12 1.24 1.73 1779.30 1.62
TABLE IX: The performance comparison between the proposed method DBgDel with three different initial remaining gene pool settings for e_coli_core. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GE gene $G_{\text{remain}}$ RC gene
Success rate 24/40 28/40 28/40 21/40
Time.Success = 1 0.59 0.68 0.60 0.59
Time.Success = 0 1.91 2.83 2.57 2.34
Time.All cases 1.12 1.33 1.12 1.17
@@
For iMM904: Table X summarizes the performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, gDel_minRN, and gMCSE in the context of targeting 700 non-extracellular space metabolites of iMM904. Regarding the success rate, DBgDel achieved success in 81 out of 700 target metabolites, meeting the minimum GR and PR threshold of 0.001 or higher during GR maximization. In comparison, GDLS failed in all 700 cases, optGene succeeded in only 27 out of 700 cases, gDel_minRN in 116 out of 700 cases, and gMCSE in 17 out of 700 cases. DBgDel demonstrated outstanding performance in average elapsed time, with a mean time of 79.92 seconds. GDLS had an average time of 243.42 seconds, optGene required significantly more extended time with an average of 1986.45 seconds, gDel_minRN had an average time of 422.32 seconds, and gMCSE had an average time of 584.24 seconds.
@@
Table XI summarizes the performance comparison between the proposed method DBgDel with three different initial remaining gene pool settings. The search space of gDel_minRN is larger than that of DBgDel, but the success rates of GE genes and $G_{remain}$ are higher. This is because GE genes and $G_{remain}$ effectively utilize information about the genes that must not be deleted, which gDel_minRN does not refer to. When utilizing Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate in achieving the minimum GR and PR threshold of 0.001 or higher during GR maximization was 81 out of 700 instances. When using the GE genes as the initial remaining gene pool, the success rate was 120 out of 700 instances. When employing $G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate remained at 126 out of 700 instances. Additionally, when employing RC genes as the initial remaining gene pool, the success rate decreased to 65 out of 700 instances. Examining the computational efficiency, the proposed method demonstrated varied performance regarding elapsed time for successful cases. For instances where success was achieved, the average elapsed time was 22.89 seconds with Predicted-$G_{\text{remain}}$ genes, 189.90 seconds with GE genes, 47.82 seconds with $G_{\text{remain}}$ genes, and 31.93 seconds with RC genes. In contrast, for cases where success was not achieved, the average elapsed time was 87.38 seconds with Predicted-$G_{\text{remain}}$ genes, 459.31 seconds with GE genes, 224.19 seconds with $G_{\text{remain}}$ genes, and 157.42 seconds with RC genes. Considering all cases, regardless of success or failure, the proposed method showed outstanding average elapsed times when using Predicted-$G_{\text{remain}}$ genes, with values of 79.92 seconds. The average elapsed times of 145.77 and 192.44 seconds remain competitive when using RC genes or $G_{\text{remain}}$ genes.
@@
TABLE X: The performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, gDel_minRN, and gMCSE in the situation of targeting non-extracellular space metabolites of iMM904. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GDLS gMCSE optGene gDel_minRN
Success rate 81/700 0/700 17/700 27/700 116/700
Avg. time 79.92 243.42 584.24 1986.45 422.32
TABLE XI: The performance comparison between the proposed method DBgDel with three different initial remaining gene pool settings for iMM904. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GE gene $G_{\text{remain}}$ RC gene
Success rate 81/700 120/700 126/700 65/700
Time.Success = 1 22.89 189.90 47.82 31.93
Time.Success = 0 87.38 459.31 224.19 157.42
Time.All cases 79.92 413.13 192.44 145.77
@@
For iML1515: Table XII summarizes the performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, gDel_minRN, and gMCSE in the context of targeting 991 non-extracellular space metabolites of iML1515. Regarding the success rate, DBgDel achieved success in 507 out of 991 target metabolites, meeting the minimum GR and PR threshold of 0.001 or higher during GR maximization. GDLS, gMCSE, and optGene failed in all 991 cases, and gDel_minRN succeeded in 508 out of 991 cases. DBgDel demonstrated outstanding performance with an average elapsed time of 431.75 seconds. GDLS had an average time of 244.58 seconds, optGene required significantly more extended time with an average of 2193.45 seconds, gDel_minRN had an average time of 2541.23 seconds, and gMCSE had an average time of 1892.39 seconds.
@@
Table XIII summarizes the performance comparison between the proposed method DBgDel with three different initial remaining gene pool settings. When utilizing Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate in achieving the minimum GR and PR threshold of 0.001 or higher during GR maximization was 507 out of 991 instances. When using the GE gene as the initial remaining gene pool, the success rate was 508 out of 991 instances. When employing $G_{\text{remain}}$ genes as the initial remaining gene pool, the success rate remained at 508 out of 991 instances. Additionally, when employing RC genes as the initial remaining gene pool, the success rate decreased to 178 out of 991 instances. Examining the computational efficiency, the proposed method demonstrated varied performance regarding elapsed time for successful cases. For instances where success was achieved, the average elapsed time was 92.46 seconds with Predicted-$G_{\text{remain}}$ genes, 481.66 seconds with GE genes, 90.30 seconds with $G_{\text{remain}}$ genes, and 80.26 seconds with RC genes. In contrast, for cases where success was not achieved, the average elapsed time was 787.17 seconds with Predicted-$G_{\text{remain}}$ genes, 4561.56 seconds with GE genes, 858.01 seconds with $G_{\text{remain}}$ genes, and 852.43 seconds with RC genes. Considering all cases, regardless of success or failure, the proposed method showed outstanding average elapsed times when using Predicted-$G_{\text{remain}}$ or $G_{\text{remain}}$ genes, with values of 431.75 and 463.70 seconds.